# Resource Allocation in STAR-RIS-Aided SWIPT with RSMA via Meta-Learning

### v2

@@ Abstract
Simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) is a cutting-edge concept for the sixth-generation (6G) wireless networks. In this letter, we propose a novel system that incorporates STAR-RIS with simultaneous wireless information and power transfer (SWIPT) using rate splitting multiple access (RSMA). The proposed system facilitates communication from a multi-antenna base station (BS) to single-antenna users in a downlink transmission. The BS concurrently sends energy and information signals to multiple energy harvesting receivers (EHRs) and information data receivers (IDRs) with the support of a deployed STAR-RIS. Furthermore, an optimization is introduced to strike a balance between users’ sum rate and the total harvested energy. To achieve this, an optimization problem is formulated to optimize the energy/information beamforming vectors at the BS, the phase shifts at the STAR-RIS, and the common message rate. Subsequently, we employ a meta deep deterministic policy gradient (Meta-DDPG) approach to solve the complex problem. Simulation results validate that the proposed algorithm significantly enhances both data rate and harvested energy in comparison to conventional DDPG.
@@ I Introduction
The concept of reconfigurable intelligent surface (RIS) is a key sixth-generation (6G) technology, attracting noteworthy attention. More specifically, each RIS consists of plenty of reflection units capable of adjusting amplitude and phase shift independently using a smart controller [1, 2]. Typically, RISs can only reflect the received incident signal and thus working in one half-plane. Therefore, a new advanced type of RIS, simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS), is recommended to overcome this limitation by enabling the full-space coverage [3]. More specifically, authors in [4] considered the deployment of STAR-RIS in a full-duplex system aiming to maximize the average weighted sum rate. Moreover, a joint deployment and beamforming design algorithm was proposed in [5] to address the sum rate maximization problem in a downlink STAR-RIS-assisted communication system. On the other hand, rate-splitting multiple access (RSMA) is a novel framework, offering substantial gains in spectral and energy efficiency. Motivated by the advantages of the integration of RIS/STAR-RIS with RSMA, a few existing works have explored the performance of this integration [6, 7, 8]. More specifically, the authors in [6] focus on maximizing the energy efficiency (EE) in a downlink RIS-aided communication system with RSMA. In [7], a proximal policy optimization method is introduced to address the sum-rate maximization problem in a downlink STAR-RIS-assisted network with RSMA. Additionally, the authors in [8] explore a two-user RIS-aided RSMA system in the context of uplink transmission, with the goal of maximizing the achievable rate. Simultaneous wireless information and power transfer (SWIPT) is a cutting-edge technology addressing energy limitations in battery-powered devices [9, 10, 11]. Furthermore, optimizing both wireless communication performance and energy transfer is crucial, leading to optimization approaches [12]. For instance, the sum rate maximization problem of multi-user SWIPT system with the deployment of an intelligent reflecting surface (IRS) is studied in [13], while the authors in [14] aim to maximize the weighted sum rate of the information users in an active IRS-assisted SWIPT system. It should be noted that this is the first work investigating a multi-user STAR-RIS-assisted SWIPT system with RSMA. In this setup, users are divided into two main groups, i.e., information decoding receivers (IDRs) and energy harvesting receivers (EHRs), to receive the signals reflected by a deployed STAR-RIS. The system performance is evaluated by jointly maximizing the total sum rate of IDRs and harvested energy by EHRs. Additionally, different from the conventional DDPG, a Meta-learning-based DDPG algorithm is formulated to solve the non-convex NP-hard problem.
@@ II System Model
where $s_{m}$ denotes the signal incident on the $m$-th element of STAR-RIS and $j$ represents the imaginary unit. Furthermore, the amplitude and phase shift of the transmission and reflection coefficients of the $m$-th element of the deployed STAR-RIS are respectively denoted by $\{\sqrt{\beta^{t}_{m}},\sqrt{\beta^{r}_{m}}\}$ and $\{\theta^{t}_{m},\theta^{r}_{m}\}$. It is worth pointing out that the amplitude and phase shift of $m$-th element of STAR-RIS in both reflection and refraction sides are within $[0,2\pi]$ and $[0,1]$, respectively. It is assumed that the reflection and transmission phase shifts $\{\theta^{t}_{m},\theta^{r}_{m}\}$ are independent, while the amplitude coefficients for both reflection and transmission are coupled. Moreover, the condition $\beta^{r}_{m}+\beta^{t}_{m}=1$ should be satisfied, as the incident signal’s energy should be equal to aggregate energy of reflected and transmitted signals, i.e., ${\left|{s_{m}^{t}}\right|^{2}}+{\left|{s_{m}^{r}}\right|^{2}}={\left|{{s_{m}}}\right|^{2}}$. With the one layer RSMA scheme at the BS, the total message of each IDR $u\in\mathcal{U}$ is divided into two main parts, i.e., a common part ${\mathbb{w}_{c}}$ and a private part $\mathbf{w}_{k}$. Then, the total transmit signal at the BS by considering both IDRs and EHRs is given by
@@
where $\mathbf{f}_{d,v}\in\mathbb{C}^{N\times 1}$ is the channel vector between the BS and $v$-th EHR, while $\mathbf{f}_{\text{S-RIS},v}\in\mathbb{C}^{M\times 1}$ denotes the channel vector between the STAR-RIS and EHR $v$. Following this, each IDR decodes its private stream after eliminating the common stream using successive interference cancellation (SIC). For the sake of simplicity, it is assumed that all IDRs can mitigate interference caused by the energy signals [13]. Hence, the received information rates of the common stream and the private stream at the $u$-th IDR can be respectively stated as follows:
@@ III Optimization Problem Formulation
In this letter, we specifically address a maximization problem for optimizing both the sum rate of IDRs and the total harvested energy at the EHRs by jointly optimizing the active beamforming vectors at the BS, i.e., $\mathbf{W}=\{\mathbf{w}_{c},\mathbf{w}_{u},\forall u\}$ and $\mathbf{P}=\{\mathbf{p}_{v},\forall v\}$ , data rate for receiving the common message $\mathbf{c}=\{C_{u},\forall u\}$, and transmission/reflection coefficients of STAR-RIS. The total sum-rate and the total harvested energy can be normalized as $R_{\text{Norm}}=\frac{\sum_{u=1}^{U}{(R_{c,u}+R_{p,u})}}{\text{max}{\{R_{c,u}+R_{p,u}\}}}$ and $P^{\textrm{Har}}_{\text{Norm}}=\frac{\sum_{v=1}^{V}P^{\textrm{Har}}_{v}}{\text{max}\{P^{\textrm{Har}}_{1},...,P^{\textrm{Har}}_{V}\}}$. Then, the maximization problem is given by
@@ IV Meta-DDPG based Learning Solution
Previous research using convex optimization techniques suggests intricate mathematical transformations for a local optimal solution. Furthermore, given the dynamic nature of wireless environments, real-time adaptability is preferred. Recent DRL methods achieve this and propose a unified solution without decomposing the problem, unlike prior convex optimization approaches. Therefore, in this letter, a Meta-DDPG approach is employed to address the optimization issue. The Meta-DDPG approach [16] is a reinforcement learning (RL) method employed due to the extensive dimensionality of both the state and action spaces. Conventional DDPG algorithm cannot quickly adapt to new and dynamic environments. Instead, Meta-learning is used to address this issue. This algorithm integrates conventional DDPG with Meta-learning to enable learning a model which rapidly adapts to a new environment. This method is based on the agent, state, action, and reward [17, 16] and utilizes actor and critic networks to learn a suitable model. In each time step $t\in[0,T]$, the agent observes a candidate state of the environment represented as $s_{t}\in\mathcal{S}$, and subsequently selects an action based on the current state $s_{t}$ which is denoted by $a_{t}\in\mathcal{A}$. After taking action $a_{t}$ in state $s_{t}$ by the agent, the environment transits to a new state $s_{t+1}\in\mathcal{S}$, in which the agent receives a numerical reward $r_{t+1}\in\mathcal{R}$. As a result, a sequence of interactions, i.e., $s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\ldots$ unfolds over time. The value of $r$ plays a crucial role as it guides the agent in selecting better actions. The policy, denoted as $\pi$, defines how the agent maps states to the actions. In the Meta-DDPG algorithm, the main purpose is to find an optimal policy denoted as $\pi^{*}$, with the ultimate aim of maximizing the reward function. To characterize the Meta-DDPG algorithm, we examine its components as follows
@@
3) Reward function: As the reward is an important part of RL, it should be designed properly to ensure the learning model, aimed at solving the optimization problem, satisfies the constraints. Therefore, the reward is defined as follows
@@ 19                         ${{\bar{\theta}}^{\pi}}_{t+1}=(1-\xi){{\bar{\theta}}^{\pi}}_{t}+\xi{\theta_{t}^{\pi}}$
where $Y=R+\tau Q(s_{t+1},\pi(s_{t+1}|\theta^{\pi})|\theta^{Q})$ and $R=\sum_{t=0}^{T}\tau^{t}r_{t}$ with $\tau\in[0,1)$. Subsequently, to update the parameters of the target critic and target actor networks, ${{\bar{\theta}}^{Q}}_{t+1}=(1-\xi){\bar{\theta}^{Q}}_{t}+{\xi{\theta_{t}^{Q}}}$ and ${{\bar{\theta}}^{\pi}}_{t+1}=(1-\xi){{\bar{\theta}}^{\pi}}_{t}+\xi{\theta_{t}^{\pi}}$ with $\xi\ll 1$ are defined. A tuple $(s_{t},a_{t},r_{t},s_{t+1})$ is stored into a buffer pool and then is updated within different episodes. The bi-level optimization problem is defined as follows:
@@ V Simulation Results
The effectiveness of the proposed method is empirically assessed with a comparison to classical methods. In the simulations, we set $M=20$, while $K$ is equal to 4. Additionally, it is supposed that the BS is equipped with $N=4$ antennas, positioned at coordinates (0,0) m. Meanwhile, the RIS is located at (100,0) m. Furthermore, the system operates within defined constraints, including $P_{\text{max}}=20~{}\textrm{dBm}$, $E_{\text{min}}=0.0004$, and $R_{\text{min}}=0.1$ bps/Hz. In Fig. 2, the Meta-DDPG depicts faster and smoother convergence compared to the DDPG approach. Moreover, combining the Meta-DDPG with STAR-RIS results in even faster convergence than with conventional RIS, highlighting the significant improvement in convergence speed. It is noteworthy that the Meta-DDPG with STAR-RIS outperforms the DDPG with STAR-RIS, emphasizing the advantages of Meta-learning with STAR-RIS.
@@ VI Complexity Analysis
Training mode
$\mathcal{O}(T\times E\times N_{t}\times S_{b}\times\sum_{i=0}^{N_{L}}\delta_{i}\delta_{i+1})$
Running mode
$\mathcal{O}(T\times E\times S_{b}\times\sum_{i=0}^{N_{L}}\delta_{i}\delta_{i+1})$
@@ VII Conclusion
In this letter, a novel approach was proposed to enhance the performance of wireless system via the deployment of an STAR-RIS. Specifically, an STAR-RIS SWIPT-assisted wireless MISO system with RSMA is considered, in which the Meta-DDPG approach is applied to solve the non-convex NP-hard problem. The users are separated into two groups, namely IDRs and EHRs which respectively receive the information and energy from the multiple-antenna BS. Furthermore, the IDRs’s sum rate maximization problem as well as the maximization of total harvested energy at the EHRs is solved by applying the Meta-DDPG algorithm.
VII-A

### v3

@@ Abstract
Simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) is a cutting-edge concept for the sixth-generation (6G) wireless networks. In this paper, we propose a novel system that incorporates STAR-RIS with simultaneous wireless information and power transfer (SWIPT) using rate splitting multiple access (RSMA). The proposed system facilitates communication from a multi-antenna base station (BS) to single-antenna users in a downlink transmission. The BS concurrently sends energy and information signals to multiple energy harvesting receivers (EHRs) and information data receivers (IDRs) with the support of a deployed STAR-RIS. Furthermore, an optimization is introduced to strike a balance between users’ sum rate and the total harvested energy. To achieve this, an optimization problem is formulated to optimize the energy/information beamforming vectors at the BS, the phase shifts at the STAR-RIS, and the common message rate. Subsequently, we employ a meta deep deterministic policy gradient (Meta-DDPG) approach to solve the complex problem. Simulation results validate that the proposed algorithm significantly enhances both data rate and harvested energy in comparison to conventional DDPG.
@@
Mojtaba Amiri Elaheh Vaezpour Sepideh Javadi Mohammad Robat Mili Halim Yanikomeroglu Mehdi Bennis
@@ I Introduction
The concept of reconfigurable intelligent surface (RIS) is a key sixth-generation (6G) technology, attracting noteworthy attention [1]. More specifically, each RIS consists of plenty of reflection units capable of adjusting amplitude and phase shift independently using a smart controller [2, 3]. Typically, RISs can only reflect the received incident signal and thus working in one half-plane. Therefore, a new advanced type of RIS, simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS), is recommended to overcome this limitation by enabling the full-space coverage [4, 5]. More specifically, the authors in [6] consider the deployment of STAR-RIS in a full-duplex system aiming to maximize the average weighted sum rate. Moreover, a joint deployment and beamforming design algorithm is proposed in [7] to address the sum rate maximization problem in a downlink STAR-RIS-assisted communication system.
On the other hand, rate-splitting multiple access (RSMA) is a novel framework, offering substantial gains in spectral and energy efficiency[8, 9, 10, 11, 12]. Motivated by the advantages of the integration of RIS/STAR-RIS with RSMA, a few existing works have explored the performance of this integration [13, 14, 15]. More specifically, the authors in [13] focus on maximizing the energy efficiency (EE) in a downlink RIS-aided communication system with RSMA. In [14], a proximal policy optimization method is introduced to address the sum-rate maximization problem in a downlink STAR-RIS-assisted network with RSMA. Additionally, the authors in [15] have explored a two-user RIS-aided RSMA system in the context of uplink transmission, with the goal of maximizing the achievable rate.
Simultaneous wireless information and power transfer (SWIPT) is a cutting-edge technology addressing energy limitations in battery-powered devices [16, 17, 18, 19]. Furthermore, optimizing both wireless communication performance and energy transfer is crucial, leading to optimization approaches [20, 21]. More specifically, the sum rate maximization problem of multi-user SWIPT system with the deployment of an intelligent reflecting surface (IRS) is studied in [22], while the authors in [23] aim to maximize the weighted sum rate of the information users in an active IRS-assisted SWIPT system. Moreover, the EE maximization problem of the active STAR-RIS-aided non-linear SWIPT system is investigated in [24], where the authors have applied both mathematical and learning methods to solve the non-convex problem. The authors in [25] aim to maximize the EE of the IRS-aided SWIPT system, where an alternating optimization (AO) algorithm is employed to solve the problem.
It should be noted that this is the first work investigating a multi-user STAR-RIS-assisted SWIPT system with RSMA. In this setup, users are divided into two main groups, i.e., information decoding receivers (IDRs) and energy harvesting receivers (EHRs), to receive the signals reflected by a deployed STAR-RIS. The system performance is evaluated by jointly maximizing the total sum rate of IDRs and harvested energy by EHRs. Additionally, different from the conventional DDPG, a Meta-learning-based DDPG algorithm is formulated to solve the non-convex NP-hard problem.
@@
The remainder of this paper is organized as follows. In Section II, the system model of a downlink STAR-RIS-assisted multiuser SWIPT system with RSMA is introduced. Section III formulates the optimization problem. In order to solve this problem, the Meta-DDPG algorithm is employed, detailed in Section IV. Furthermore, simulation results are expressed in Section V, while complexity analysis is shown in Section VI. Finally, Section VII concludes the paper.
Notation: In this paper, boldface capital letters and lowercase letters represent matrices and vectors, respectively. $\mathbb{E}(.)$ and (.$(.)^{H}$ denote the expectation value and conjugate transposition, respectively, while $\mathbb{C}$ denotes the complex matrix.
@@ II System Model
where $s_{m}$ denotes the signal incident on the $m$-th element of STAR-RIS and $j$ represents the imaginary unit. Furthermore, the amplitude and phase shift of the transmission and reflection coefficients of the $m$-th element of the deployed STAR-RIS are respectively denoted by $\{\sqrt{\beta^{t}_{m}},\sqrt{\beta^{r}_{m}}\}$ and $\{\theta^{t}_{m},\theta^{r}_{m}\}$. It is worth pointing out that the amplitude and phase shift of $m$-th element of STAR-RIS in both reflection and refraction sides are within $[0,2\pi]$ and $[0,1]$, respectively. It is assumed that the reflection and transmission phase shifts $\{\theta^{t}_{m},\theta^{r}_{m}\}$ are independent, while the amplitude coefficients for both reflection and transmission are coupled. Moreover, the condition $\beta^{r}_{m}+\beta^{t}_{m}=1$ should be satisfied, as the incident signal’s energy should be equal to aggregate energy of reflected and transmitted signals, i.e., ${\left|{s_{m}^{t}}\right|^{2}}+{\left|{s_{m}^{r}}\right|^{2}}={\left|{{s_{m}}}\right|^{2}}$.
With the one layer RSMA scheme at the BS, the total message of each IDR $u\in\mathcal{U}$ is divided into two main parts, i.e., a common part ${\mathbb{w}_{c}}$ and a private part $\mathbf{w}_{k}$. Then, the total transmit signal at the BS by considering both IDRs and EHRs is given by
@@
where $\mathbf{f}_{d,v}\in\mathbb{C}^{N\times 1}$ is the channel vector between the BS and $v$-th EHR, while $\mathbf{f}_{\text{S-RIS},v}\in\mathbb{C}^{M\times 1}$ denotes the channel vector between the STAR-RIS and EHR $v$.
Following this, each IDR decodes its private stream after eliminating the common stream using successive interference cancellation (SIC). For the sake of simplicity, it is assumed that all IDRs can mitigate interference caused by the energy signals [22]. Hence, the received information rates of the common stream and the private stream at the $u$-th IDR can be respectively stated as follows [11, 12]:
@@ III Optimization Problem Formulation
In this paper, we specifically address a maximization problem for optimizing both the sum rate of IDRs and the total harvested energy at the EHRs by jointly optimizing the active beamforming vectors at the BS, i.e., $\mathbf{W}=\{\mathbf{w}_{c},\mathbf{w}_{u},\forall u\}$ and $\mathbf{P}=\{\mathbf{p}_{v},\forall v\}$, data rate for receiving the common message $\mathbf{c}=\{C_{u},\forall u\}$, and transmission/reflection coefficients of STAR-RIS. The total sum-rate and the total harvested energy can be normalized as $R_{\text{Norm}}=\frac{\sum_{u=1}^{U}{(R_{c,u}+R_{p,u})}}{\text{max}{\{R_{c,u}+R_{p,u}\}}}$ and $P^{\textrm{Har}}_{\text{Norm}}=\frac{\sum_{v=1}^{V}P^{\textrm{Har}}_{v}}{\text{max}\{P^{\textrm{Har}}_{1},...,P^{\textrm{Har}}_{V}\}}$. Then, the maximization problem is given by
@@ IV Meta-DDPG based Learning Solution
Previous research using convex optimization techniques suggests intricate mathematical transformations for a local optimal solution. Furthermore, given the dynamic nature of wireless environments, real-time adaptability is preferred. Recent DRL methods achieve this and propose a unified solution without decomposing the problem, unlike prior convex optimization approaches. Therefore, in this paper, a Meta-DDPG approach is employed to address the optimization issue. The Meta-DDPG approach [27] is a reinforcement learning (RL) method employed due to the extensive dimensionality of both the state and action spaces. Conventional DDPG algorithm cannot quickly adapt to new and dynamic environments. Instead, Meta-learning is used to address this issue. This algorithm integrates conventional DDPG with Meta-learning to enable learning a model which rapidly adapts to a new environment. This method is based on the agent, state, action, and reward [28, 27, 29] and utilizes actor and critic networks to learn a suitable model.
In each time step $t\in[0,T]$, the agent observes a candidate state of the environment represented as $s_{t}\in\mathcal{S}$, and subsequently selects an action based on the current state $s_{t}$ which is denoted by $a_{t}\in\mathcal{A}$. After taking action $a_{t}$ in state $s_{t}$ by the agent, the environment transits to a new state $s_{t+1}\in\mathcal{S}$, in which the agent receives a numerical reward $r_{t+1}\in\mathcal{R}$. As a result, a sequence of interactions, i.e., $s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\ldots$, unfolds over time. The value of $r$ plays a crucial role as it guides the agent in selecting better actions. The policy, denoted as $\pi$, defines how the agent maps states to the actions. In the Meta-DDPG algorithm, the main purpose is to find an optimal policy denoted as $\pi^{*}$, with the ultimate aim of maximizing the reward function. To characterize the Meta-DDPG algorithm, we examine its components as follows:
@@
3) Reward function: As the reward is an important part of RL, it should be designed properly to ensure the learning model, aimed at solving the optimization problem, satisfies the constraints. Therefore, the reward is defined as follows:
@@ 19                         ${{\bar{\theta}}^{\pi}}_{t+1}=(1-\xi){{\bar{\theta}}^{\pi}}_{t}+\xi{\theta_{t}^{\pi}}$
where $Y=R+\tau Q(s_{t+1},\pi(s_{t+1}|\theta^{\pi})|\theta^{Q})$ and $R=\sum_{t=0}^{T}\tau^{t}r_{t}$ with $\tau\in[0,1)$. Subsequently, to update the parameters of the target critic and target actor networks, ${{\bar{\theta}}^{Q}}_{t+1}=(1-\xi){\bar{\theta}^{Q}}_{t}+{\xi{\theta_{t}^{Q}}}$ and ${{\bar{\theta}}^{\pi}}_{t+1}=(1-\xi){{\bar{\theta}}^{\pi}}_{t}+\xi{\theta_{t}^{\pi}}$ with $\xi\ll 1$ are defined.
A tuple $(s_{t},a_{t},r_{t},s_{t+1})$ is stored into a buffer pool and then is updated within different episodes. The bi-level optimization problem is defined as follows:
@@ V Simulation Results
The effectiveness of the proposed method is empirically assessed with a comparison to classical methods. In the simulations, we set $M=20$, while $K$ is equal to 4. Additionally, it is supposed that the BS is equipped with $N=4$ antennas, positioned at coordinates (0,0) m. Meanwhile, the RIS is located at (100,0) m. Furthermore, the system operates within defined constraints, including $P_{\text{max}}=20~{}\textrm{dBm}$, $E_{\text{min}}=0.0004$, and $R_{\text{min}}=0.1$ bps/Hz.
In Fig. 2, the reward for various episodes is shown, where the Meta-DDPG depicts faster and smoother convergence compared to the DDPG approach. Moreover, combining the Meta-DDPG with STAR-RIS results in even faster convergence than with conventional RIS, highlighting the significant improvement in convergence speed. It is noteworthy that the Meta-DDPG with STAR-RIS outperforms the DDPG with STAR-RIS, emphasizing the advantages of Meta-learning with STAR-RIS.
@@
In the following, a simulation is conducted to assess the ANHE across a range of different values for $\alpha$, as depicted in Fig. 5. This figure reinforces the point that the integration of the Meta-DDPG approach with STAR-RIS yields acceptable results and enhances energy harvesting performance.
Figure 5: The ANHE versus $\alpha$.
@@ VI Complexity Analysis
Training mode $\mathcal{O}(T\times E\times N_{t}\times S_{b}\times\sum_{i=0}^{N_{L}}\delta_{i}\delta_{i+1})$
Running mode $\mathcal{O}(T\times E\times S_{b}\times\sum_{i=0}^{N_{L}}\delta_{i}\delta_{i+1})$
@@ VII Conclusion
In this paper, a novel approach was proposed to enhance the performance of wireless system via the deployment of an STAR-RIS. More specifically, an STAR-RIS SWIPT-assisted wireless MISO system with RSMA is considered, in which the Meta-DDPG approach is applied to solve the non-convex NP-hard problem. In this setup, the users are separated into two groups, namely IDRs and EHRs which receive the information and energy from the multiple-antenna BS, respectively. Furthermore, the IDRs’s sum rate maximization problem as well as the maximization of total harvested energy at the EHRs is solved by applying the Meta-DDPG algorithm. Simulation results confirmed the superiority of the proposed Meta-DDPG algorithm in both data rate and harvested energy enhancements compared to the conventional DDPG.